OpenStudy (anonymous):

Find all pears of (a,b) satisfy : a^4 - 49a^2 + 1 = b^2. (a,b are natural number)

OpenStudy (anonymous):

i think one of solution is (7,1), but i got the answer using trial and error method :) what about u @mukushla ??

OpenStudy (anonymous):

im workin on it... one method is making a quadratic equation in terms of \(a^2\) ...then the Discriminant of that quadratic must be perfect square...let me check it

OpenStudy (anonymous):

let \(a^2=m\) then u have\[m^2 - 49m + 1 - b^2=0\]Discriminant= \(49^2-4(1-b^2)=n^2\) must be a perfect square it gives\[(n-2b)(n+2b)=2397=3*17*47\] now there are few options : 1............\(n-2b=1\) and \(n+2b=2397\) 2............\(n-2b=3\) and \(n+2b=799\) 3............\(n-2b=17\) and \(n+2b=141\) 4............\(n-2b=47\) and \(n+2b=51\)

OpenStudy (anonymous):

Why discriminant of that quadratic must be perfect square??

OpenStudy (anonymous):

ohh i see, so that the roots of quadtic equation be integer, isn't ?

OpenStudy (anonymous):

lets solve it completely...

OpenStudy (anonymous):

i late get information from u mukushla... :D So, if i solve the last that question. I got (n,b) = (1199,599), (401, 199), (79,31), and (49,1) what next i do??

OpenStudy (anonymous):

now we must plug values of \(b\) in the quadratic and solve for \(m\) .... \(m\) must be a perfect square too because we solved equation with \(m=a^2\)

OpenStudy (anonymous):

solve and tell me what u get?

OpenStudy (anonymous):

only m?

OpenStudy (anonymous):

\(b\) gives \(m\) and \(m\) gives \(a\)

OpenStudy (anonymous):

not conencted in my mind... (b gives m and m gives a, what mean is it)

OpenStudy (anonymous):

subtitute for b to m^2−49m+1−b^2=0 ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

if b=1, i got m^2-49m=0 => m(m-49)=0 m=0 (not satisfy) m=49 from a^2 = m a^2 = 49, so a=7 (-7 not natural number) so the first pear for (a,b) = (7,1) and next next next .... right mukushla?

OpenStudy (anonymous):

yep....:D

OpenStudy (anonymous):

i was just solved it, hehe... thanks million for u mukushla...

OpenStudy (anonymous):

You are Welcome

OpenStudy (anonymous):

but im not sure i can solve again if given like this problems... :D

OpenStudy (anonymous):

im sure u can...

OpenStudy (anonymous):

if the problems same like this, hehe but i will study hard to get it......